# F-BIAS team

### v1

@@ Abstract
We discuss in this article the creation and organization of a national core facility for bioimage analysis, based on a distributed team. F-BIAS federates bioimage analysts across France and relies on them to deliver services to the researchers of this territory. The main challenge in implementing this structure is to provide significant scientific value to its members, thereby encouraging their active participation and persuading their respective host teams to support their involvement. F-BIAS accomplished this by creating a professional network that mitigates the negative effects of isolation experienced by its members, who are often the sole bioimage analyst within their local teams, and fosters the development of their technical skills. In a second phase we capitalized on F-BIAS to create a virtual, remotely-operating core facility for bioimage analysis, offering consultations and collaborative project services to the scientific community of France. The services are organized so that they also contribute to the technical proficiency of the analysts. To promote the creation of similar structures, we present and discuss here the organization of this nationally distributed bioimage analysis service core, highlighting successes and challenges.
@@
Key points We describe our experience in creating a nationwide core facility offering services in bioimage analysis. This virtual facility federates existing resources scattered across the territory and within different research institutions. It attracted and retained analysts thanks to the scientific added-value it brought to them. The distributed core facility efficiently addresses the needs of imaging-based research projects, in particular for researchers previously without access to image analysis expertise. It also mitigates the risk of isolation for analyst members that are often the sole expert in image analysis in their local teams. We identify and share the critical components necessary for the success of similar endeavors, to facilitate reproducibility.
@@ Introduction
Core facilities are today a well-established part of the research landscape in academia. They are run by dedicated staff and provide expertise and assistance to research projects in biology (17). While centralizing computational resources for bioinformatics is well established (18), bioimage analysis core facilities - dedicated structures that house analysts and provide image analysis services to a broad scientific community (16, 19) - are only emerging. Most of the time analysts are embedded in large microscopy cores. This has some clear advantages in a holistic approach: bioimage analysis is indissociable from imaging. For instance, a challenging bioimage analysis project can be made simpler by choosing the right imaging modality.
@@
France-BioImaging (FBI), the French node of EuroBioImaging (27), is the French nationwide infrastructure for biological imaging (28). It is a distributed infrastructure that provides its users with quick access to cutting-edge and innovative microscopy, associated expertise, labeling methods, image analysis, and tailored training. It is composed of microscopy core facilities but also of research and development teams in this different fields in order to foster technological transfer of innovation to users. Many of the analysts working in French core facilities followed the isolated analyst model described above (Figure 1). This led us to build a distributed structure, supported by FBI, to mitigate the issues of isolation on the scale of a country. In turn, it opened up an opportunity to build a nationwide, virtual, remotely-operating core facility for bioimage analysis. In the following we report the creation of F-BIAS: France-BioImaging Analysts the federation of analysts working in French core facilities. F-BIAS was initially created to have a strong scientific value to analysts. In a second phase, it was used to offer bioimage analysis services to the French scientific community as a distributed core. We describe below how F-BIAS was designed, what rules were used and what can be achieved with such an organization. We also discuss its limitations and challenges and share our experience, hoping that it will be useful for other bioimage analysis communities over the world.
@@
F-BIAS was created to federate analysts scattered across France in order to provide an adequate professional environment and to address the needs listed in the previous section. To this aim, a few rules were defined. First, members must be bioimage analysts working in a core facility in France whose main mission is to deliver bioimage analysis services to the wide scientific community. Image processing scientists working in research labs cannot join F-BIAS as support is not part of their mission and their work mainly focuses on their own projects. Second, members have to contribute to the services of the nationwide core-facility described below, mainly by participating in the open-desk sessions. Third, meetings and exchanges are private and confidential. Analysts are asked not to disclose any information discussed during meetings unless agreed upon and to immediately voice concerns if they see a possible conflict of interest when a topic is brought up by a peer. This helps to create a safe discussion place where confidentiality issues are addressed on a case-by-case basis.
@@
In practice, F-BIAS was initiated mid 2021 by inviting analysts working in FBI core facilities to join under the rules and scope outlined above. The resulting network started with 10 members and has since expanded to include 20 members today (Figure 1). This turns out to be a manageable size, handy to maintain focused, lively, and efficient exchanges. All members have the same line of work with a common commitment to support a scientific community with bioimage analysis, but they show a diversity of backgrounds (applied mathematics, biology, computer science, physics), domains of applications (cell biology, developmental biology, histopathology, microbiology, plant biology, …) and career stage. The initial activities involved establishing monthly meetings and deploying team chat tools to favor instantaneous exchange and structured communication. Meetings are prepared in advance to include 2-5 topics including 1/ sharing preview versions of tools and pipelines developed by members and collecting feedback, 2/ requesting help by peers on a user project, 3/ discussing an article or a new bioimage analysis tool. Additional yearly in-person hackathons of 2-3 days and other gatherings during FBI meetings help to maintain team spirit. F-BIAS governance is horizontal and relies on two to three members for event organization, service coordination and reporting to FBI, our umbrella infrastructure.
@@
F-BIAS aims to provide bioimage analysis services to the entire scientific community in France, particularly life scientists with no analyst onsite. To fulfill this mission, F-BIAS offers bioimage analysis consultations to researchers in France in the form of an open-desk, organized every other month, and communicated through FBI newsletter and website. For each open-desk, two to six online sessions of 1 hour are planned, with users booking one of the sessions in advance. Users are required to upload images before the session with a description of their needs. Descriptions are then used to select 2 to 3 F-BIAS members that will join the user in an online session of 1 hour (Figure 2A). Analysts can offer advice on which tool to use, how to use it with online demonstration on the user data, implement a preliminary version of an analysis pipeline, or create small macros for analysis or task automation. Within the allocated time of the session, the goal is to address the user’s specific question and analysts are required to prototype the solutions and advice they provide. Most of the time, the different analysts propose different approaches. Consultations are free and users are expected to acknowledge the FBI infrastructure in their publications.
@@
Consultations are best suited when the image analysis task can be solved with existing tools or well-established pipelines. However, imaging-based research feeds on innovation. In some cases, researchers have questions that have never been addressed before and for which no tools are suited. A significant part of the analysts’ scientific added-value is their ability to adapt existing tools and to create bespoke image analysis pipelines, in the framework of collaborative projects. In the scope of bioimage analysis, this term relates to the elaboration and delivery of an image analysis pipeline or of image analysis results by the analyst to the user, with both parties closely working together. The project corresponds to an articulated request from the user but involves an original contribution and some scientific responsibility from the analyst. F-BIAS aims at offering such services at a national level.
@@
While consultations require moderate time commitments, collaborative projects take significant time during which an analyst becomes unavailable. Understandably, both analysts and their supervisors expressed caution when joining the initiative. An important factor in overcoming these reservations was a proof-of-concept collaborative project established in 2023, involving an analyst in Nantes, developing an ImageJ macro for a user in Montpellier, and tutored by another analyst in Strasbourg. This demonstrated the feasibility of delivering high-quality services remotely and integrating them into the existing service offers of host core facilities. This success facilitated the allocation of two junior engineer positions dedicated to F-BIAS supported by the FBI umbrella infrastructure. Thanks to these recruitments, within the first six months of 2024, we were able to accept and plan eight requests for collaborative projects. Out of these eight requests, three were completed and one was deemed unfeasible. User requests cover a wide range of domains in biology and image analysis (Table 1), which highlights the diversity of expertise F-BIAS can offer, with a distributed team coming from diverse backgrounds.
@@
Users can request a collaborative project by direct contact, via Euro-BioImaging user access website or during an open-desk session. The request is then discussed among the F-BIAS team - that serves as a selection committee - to evaluate the project’s feasibility. If deemed feasible, one analyst with the required expertise and availability takes over the project. It evolves iteratively with regular discussions between the biologist and the analyst. Additionally, the project’s analyst may be paired with another analyst from a different site providing technical mentorship. This again results in capitalizing on user services to improve the skills and expertise of analysts (Figure 2B).
@@
Figure 2: Overview of the bioimage analysis services provided by F-BIAS. A. The consultations services take the shape of open-desk sessions of one hour organized regularly. A user can book such a session in an online form, providing a description of the image analysis problem they want to discuss, and uploading a few example images. During the preparation phase, the requests are each allocated to an open-desk session, with a panel of 2 to 3 analysts selected according to their expertise. All the sessions happen the same day, with 2 to 3 running in parallel. The analysts join the user in an online meeting of 1 hour and work on the user’s problem during this time, actively prototyping the solutions they propose. The output of a session can be advice on how to address the problem, small analysis pipeline or small macros. B. When a question cannot be addressed in one hour and / or requires some development, a user can request a collaborative project, during an open-desk session or by direct contact. Project requests are then evaluated by the F-BIAS team, which estimates the project cost, start delay, and duration. Upon acceptance, an analyst is assigned based on availability and expertise. As a potential opportunity. a senior analyst from another site will provide mentoring to the project’s analyst. The user and analyst collaborate asynchronously, with regular analyst reports, using various communication media (emails, online/onsite meetings). Project outputs may include analysis pipelines, results, and reports.
@@
The added-value of a distributed image analysis core-facility
@@
F-BIAS successfully addresses isolation issues mentioned in the above. Within F-BIAS, analysts exchange knowledge and good practices, teach each other analysis techniques and share information on key publications, online resources or conferences. Experienced members can also mentor early-career peers and help them navigate the global bioimage analysis world. Conversely, early-career analysts challenge the techniques senior analysts are used to relying on, and help them renew their technical portfolio. This is particularly valuable given the young age of the field in academia. The consultations with several analysts amounted to an opportunity for learning how others approach a new problem and exchanging ideas. It also provided the group’s developers with a chance to demonstrate new tools and developments, have them tested by other analysts, and receive feedback. F-BIAS is highly valued by its members as a vital platform for honing skills, exploring new techniques, and discovering new tools.
@@
The scientific community of France is the main target of this endeavor, especially biologists who do not have access to bioimage analysis expertise onsite, the most common situation in the country. These scientists now have access to the expertise of about 20 analysts and the means to leverage it for their projects. Consultations are tailored for efficiency and maximizing the scientific output of a one-hour session. F-BIAS analysts master a wide range of image analysis languages and packages from scripting to advanced analysis modules and AI libraries. By describing their needs in the online registration form, users will be put in contact with analysts that have the given expertise to answer their question, when possible. From the user viewpoint, this process is much simpler than having to identify the required expertise to find the right facility to contact. This one-hour consulting service with several experts is free of charge and requires little effort beside registration and sending test data beforehand. Finally, when a question cannot be solved in one hour and the user does not have resources to develop a suggested solution, F-BIAS can initiate a collaborative project with the user. F-BIAS truly succeeds in overcoming the technological barriers associated with bioimage analysis for researchers in the France territory.
@@
The integration of F-BIAS within France-BioImaging enables it to offer an efficient service in image analysis that complements other services. Moreover, F-BIAS reached critical size to foster beneficial interactions with other projects of France-BioImaging related to bioimage analysis. For instance, the FBI.Data project is focused on scientific image data management, offering tools for intelligent data archival, high-performance computing, and implementation of FAIR data principles. FBI.Data can capitalize on F-BIAS to test and deploy its solutions in host core facilities, and reach out to end users, using F-BIAS members as a medium. FBI also supports the development of large scientific framework for image analysis such as Icy (29), BioImageIT (30) and MorphoNet (31). F-BIAS members serve as early adopters, beta testers and feedback source for these projects, offering a wide range of biological applications to disseminate the software through the F-BIAS user base. Finally, F-BIAS can redirect projects that fall outside the scope of its services to research groups of FBI focused on image processing.
@@
The most significant challenge is to secure human resources, in particular for the collaborative project services. Indeed, a major issue in operating the group and bioimage analysis core facilities in general is the limited number of dedicated analysts compared to the needs of the biological community. F-BIAS is a virtual structure, initially built on existing resources that are federated from host core facilities. At its creation, all members were already saturated or close to being saturated with requests from their local scientific community. This low availability is the main factor delaying image analysis projects (18). The inception of F-BIAS was facilitated by the scientific value it offered to members and the minimal time commitment required for participation in consultations and exchanges. This helped in building and maintaining a sufficiently large team of willing analysts. Additionally, the positions granted by FBI for collaborative projects further strengthened our services.
@@
Ensuring the scientific added-value of internal meetings
The scientific added-value analysts find in F-BIAS is crucial for the success of the initiative. It helped convince the analysts’ managers to allocate time to the initiative, and helped retain analysts and prevent dropouts. It is therefore important to control the content of F-BIAS regular and annual meetings. We do this by enforcing the rules described above and focus on technical exchanges. Regular meetings are mainly focused on scientific content, asking for each other’s help or showing a new technique, and round tables on more organizational aspects are reserved for our annual hackathons.
@@
The desirable components a distributed image analysis core facility
There are a few other initiatives building nationwide services for bioimage analysis. Among them we can mention the SciLifeLab BioImage Informatics infrastructure in Sweden (33) and the National Facility for Data Handling & Analysis in Italy (34). They have the advantage of being centralized in one site and enjoying human resources dedicated to these projects from their inception. On our side we started as a virtual structure, built from existing resources initially with other duties, spread over the country. Our approach described above managed to circumvent these challenges and even converted them into opportunities. Nonetheless we identified a few key components, additional to the human resources and online tools discussed above, that would facilitate reproducing such a service elsewhere. As we did not benefit from all of them immediately, we can attest to their importance.

### v2

@@ Abstract
We discuss in this article the creation and organization of a national core facility for bioimage analysis, based on a distributed team. F-BIAS federates bioimage analysts across France and relies on them to deliver services to the researchers of this territory. The main challenge in implementing this structure is to ensure significant scientific value to the analysts, thereby encouraging their active participation and persuading their respective host teams to support their involvement. F-BIAS accomplished this by creating a professional network that mitigates the negative effects of isolation experienced by its members, who are often the sole bioimage analyst within their local teams, and fosters the development of their technical skills. In a second phase we capitalized on F-BIAS to create a virtual, remotely-operating core facility for bioimage analysis, offering consultations and collaborative project services to the scientific community of France. The services are organized so that they also contribute to the technical proficiency of the analysts. To promote the creation of similar structures, we present and discuss here the organization of this nationally distributed bioimage analysis service core, highlighting successes and challenges.
@@
Key points We describe our experience in creating a nationwide core facility offering services in bioimage analysis. This virtual facility federates existing resources scattered across the territory and within different research institutions. It attracted and retained analysts thanks to the scientific added value it brought to them. The distributed core facility efficiently addresses the needs of imaging-based research projects, in particular for researchers previously without access to image analysis expertise. It also mitigates the risk of isolation for analyst members that are often the sole expert in image analysis in their local teams. We identify and share the critical components necessary for the success of similar endeavors, to facilitate reproducibility.
@@
{acronyms}
CNRS: Centre national de la recherche scientifique. GloBIAS: Global BioImage Analysts’ Society. INSERM: Institut national de la santé et de la recherche médicale. NEUBIAS: Network of European BioImage Analysts. RMS: Royal Microscopical Society. RTmfm: Réseau Technologique Microscopie Fluorescence Multidimensionnelle.
@@ Introduction
Core facilities are today a well-established part of the research landscape in academia. They are run by dedicated staff and provide expertise and assistance to research projects in biology (17). While centralizing computational resources for bioinformatics is well established (18), bioimage analysis core facilities - dedicated structures that house analysts and provide image analysis services to a broad scientific community (16, 19) - are only emerging. Most of the time analysts, when they are present, are embedded in large microscopy cores. This has some clear advantages in a holistic approach: bioimage analysis is indissociable from imaging. For instance, a challenging bioimage analysis project can be made simpler by choosing the right imaging modality.
@@
France-BioImaging (FBI), the French node of EuroBioImaging (27), is the French nationwide infrastructure for biological imaging (28). It is a distributed infrastructure that provides its users with quick access to cutting-edge and innovative microscopy, associated expertise, labeling methods, image analysis, and tailored training. It is composed of microscopy core facilities but also of research and development teams in these different fields in order to foster technological transfer of innovation to users. Many of the analysts working in French core facilities followed the isolated analyst model described above (Figure 1). This led us to build a distributed structure, supported by FBI, to mitigate the issues of isolation on the scale of a country. In turn, it opened up an opportunity to build a nationwide, virtual, remotely-operating core facility for bioimage analysis. In the following we report the creation of F-BIAS: France-BioImaging Analysts the federation of analysts working in French core facilities. F-BIAS was initially created to have a strong scientific value to analysts. In a second phase, it was used to offer bioimage analysis services to the French scientific community as a distributed core. We describe below how F-BIAS was designed, what rules were used and what can be achieved with such an organization. We also discuss its limitations and challenges and share our experience, hoping that it will be useful for other bioimage analysis communities over the world.
@@
While developing F-BIAS, we encountered several challenges throughout the different phases of deployment, as outlined in Table 1 and detailed below.
{tblr}
width = colspec = Q[25]Q[50]Q[115], cells = m, fg=Tundora, rowodd = AthensGray, row1 = Nandor,fg=white, column1 = l, column2 = l, column3 = l, hlines, vlines = Black, vline1 = -LunarGreen, Steps & Challenges Approach
[…]
[r=3]cInception of the F-BIAS network Defining a target community Restricted to research engineers working in a FBI core facility, providing services to the scientific community and whose main occupation is bioimage analysis.
Defining requirements for membership Regular attendance at monthly meetings and open desks is required.
Recruiting initial members Personal 1-1 calls with tentative members, advocating the network and collecting expectations, focused on the needs of isolated analysts.
[…]
[r=3]cMonthly meetings organization Planning the meeting sessions Poll at the beginning of the year to define the optimal meeting time, and sending invitations.
Establishing rules for exchange of information Schedule private meetings (without users) and secure a commitment from attendees to disclose any conflict of interest related to projects presented during the meeting.
Encouraging attendance Facilitate continuous discussions on the instant chat system between sessions, and send a reminder one week before each meeting. Restrict meetings content to analysts’ occupation: get advice and help on user projects, discuss new tools or techniques developed by a member or published by others.
[…]
[r=2]cYearly meetings organization Coordinating logistics Hosted and organized by an analyst member in the core facility they belong to.
Defining the meeting scientific content Use the instant chat system for discussions to define the program, which will focus on technical talks for upskilling and general discussions on organizing the bioimage analysis core facility’s work.
[…]
[r=5]cOpen desk sessions (Figure 2A) Planning the next session Fixed during monthly meetings to have as many available analysts as possible. One open-desk session can last 1-3 hours, have 1-6 slots of one hour, with possibly 2 slots running in parallel.
Communicating about the session Announcement to all FBI users by the FBI External Affairs Manager & creation of an online page on FBI website for enrollment.
Assembling the analysts group for one slot Analysts are mixed as much as possible based on their profiles to achieve the following: a) encourage them to learn and explore different approaches and tools, b) provide users with a broader range of ideas and perspectives.
Maximizing the efficiency and the scientific value of the open desk slot Users are requested to send example data in advance so that analysts can work on them.
Keeping track of the session content Shared Google Doc file with user information as well as the proposed solution, reasonably detailed.
[…]
[r=7]cCollaborative projects (Figure 2B) Obtaining dedicated human resources Advocate for the need for these resources within the FBI umbrella by showcasing the success of a proof-of-concept collaborative project.
Defining a recharge policy Discussions with the FBI executive board to define an hourly fee that can be subsidized by FBI. Discussions with members core facilities to implement the subsidized fee to F-BIAS collaborative projects.
Defining an entry point for user requests Discussions with a user during an open desk session or via direct contact. The project request is then formalized by the user through the Euro-BioImaging Access Portal.
Choosing the analysts for collaboration Discussion during monthly meeting to a) check availability and expertise of analysts, b) get some advices on the project and c) potentially elect a mentor analyst that will support the collaborating analyst.
Performing follow-up for the collaborative project Regular meetings are held between the analyst and the user, the collaborating analyst and the mentor analyst, or all parties together. The collaborative analyst can also rely on the F-BIAS group. Project management will use the tools and policies of the analyst’s core facility.
Sharing project data Perspective: use of an OMERO database instance deployed with the help of FBI.Data project across the different member structures.
Charging the user for the project Billing is managed by the host core facility of the collaborating analyst, based on the agreed-upon fees.
Table 1: List of the milestones, their challenges, and the approaches we used to address them.
@@
In France, research support positions are typically held by technicians or research engineers, who have a dedicated career path distinct from that of researchers. In core facilities, these positions can be either fixed-term contracts or permanent, and they are often funded by the institutions that host the core facilities (such as universities, Institut Pasteur, Institut Curie, etc.) or by research organizations (like CNRS, INSERM, etc.) that fund these cores. The mission of analysts in these positions is to support their local scientific community through the activities of the core facility in which they are embedded.
@@
F-BIAS was created to federate analysts scattered across France in order to provide an adequate professional environment and to address the needs listed in the previous section. To this aim, a few rules were defined. First, members must be bioimage analysts working in a core facility in France whose main mission is to deliver bioimage analysis services to the wide scientific community. Image processing scientists working in research labs cannot join F-BIAS as support is not part of their mission and their work mainly focuses on their own projects. Second, members have to contribute to the services of the nationwide core facility described below, mainly by participating in the open desk sessions. Third, meetings and exchanges are private and confidential. Analysts are asked not to disclose any information discussed during meetings unless agreed upon and to immediately voice concerns if they see a possible conflict of interest when a topic is brought up by a peer. This helps to create a safe discussion place where confidentiality issues are addressed on a case-by-case basis.
@@
In practice, F-BIAS was initiated mid 2021 by inviting analysts working in FBI core facilities to join under the rules and scope outlined above. As candidate members already held research support positions, no initial funding was required. The initial effort involved convincing the analysts and their managers to allocate some of the analysts’ time for this network. Since some of the research organizations in France are national entities (such as CNRS and INSERM), they create incentives to participate in national initiatives like F-BIAS. As such, F-BIAS started as virtual network, federating existing resources. The resulting network started with 10 members and has since expanded to include 20 members today (Figure 1). This turns out to be a manageable size, handy to maintain focused, lively, and efficient exchanges. All members have the same line of work with a common commitment to support a scientific community with bioimage analysis, but they show a diversity of backgrounds (applied mathematics, biology, computer science, physics), domains of applications (cell biology, developmental biology, histopathology, microbiology, plant biology, …) and career stage. The initial activities involved establishing monthly meetings and deploying team chat tools to favor instantaneous exchange and structured communication. Meetings are prepared in advance to include 2-5 topics including 1/ sharing preview versions of tools and pipelines developed by members and collecting feedback, 2/ requesting help by peers on a user project, 3/ discussing an article or a new bioimage analysis tool. Approximately 30 meetings have taken place since the creation of F-BIAS. Additional yearly in-person of 2-3 days and other gatherings during FBI meetings help to maintain team spirit. Most of the analysts are also active in other networks such as GloBIAS or CZI. When a member wishes to learn a new technique, we inquire within the group to see if someone is knowledgeable in that area and willing to conduct a short demonstration or workshop at our annual or monthly meetings. F-BIAS governance is horizontal and relies on two to three members for event organization, service coordination and reporting to FBI, our umbrella infrastructure.
@@
F-BIAS aims to provide bioimage analysis services to the entire scientific community in France, particularly life scientists with no analyst onsite. To fulfill this mission, F-BIAS offers bioimage analysis consultations to researchers in France in the form of an open desk, organized every other month, and communicated through FBI newsletter and website. For each open desk, two to six online sessions of 1 hour are planned, with users booking one of the sessions in advance. Users are required to upload images before the session with a description of their needs. Descriptions are then used to select 2 to 3 F-BIAS members that will join the user in an online session of 1 hour (Figure 2A). Analysts can offer advice on which tool to use, how to use it with online demonstration on the user data, implement a preliminary version of an analysis pipeline, or create small macros for analysis or task automation. Within the allocated time of the session, the goal is to address the user’s specific question and analysts are required to prototype the solutions and advice they provide. Most of the time, the different analysts propose different approaches. Consultations are free and users are expected to acknowledge the FBI infrastructure in their publications.
@@
A shared document is completed at the end of each session, summarizing the attending analysts and the solution proposed to the user as well as any raised issues (such as the need for better resolution or advices for improved imaging etc). This ensures the continuity if the user attends another open desk with different analysts.
@@
Consultations are best suited when the image analysis task can be solved with existing tools or well-established pipelines. However, imaging-based research feeds on innovation. In some cases, researchers have questions that have never been addressed before and for which no tools are suited. A significant part of the analysts’ scientific added value is their ability to adapt existing tools and to create bespoke image analysis pipelines, in the framework of collaborative projects. In the scope of bioimage analysis, this term relates to the elaboration and delivery of an image analysis pipeline or of image analysis results by the analyst to the user, with both parties closely working together. The project corresponds to an articulated request from the user but involves an original contribution and some scientific responsibility from the analyst. F-BIAS aims at offering such services at a national level.
@@
While consultations require moderate time commitments, collaborative projects require a significant amount of the analyst time. Understandably, both analysts and their supervisors expressed caution when joining the initiative. The main mission of an analyst is to support the users of their local core facility. We observed that all the initial members of F-BIAS were already fully occupied with local projects, which naturally limited their ability to accept new projects from outside their hosting core facilities. Consequently, we turned to FBI, acting as the national infrastructure umbrella, to provide dedicated resources. An important factor in advocating for these resources was a proof-of-concept collaborative project established in 2023, involving an analyst, a user and a mentoring analyst in three distant cities. It demonstrated the feasibility of delivering high-quality services remotely and integrating them into the existing service offers of host core facilities while sharing the work load and allowing mentoring between analysts. This success facilitated the allocation of two junior engineer positions dedicated to F-BIAS projects, financially supported by the FBI umbrella infrastructure, and that were embedded in two cores of F-BIAS. Thanks to these recruitments, within the first six months of 2024, we were able to accept and plan eight requests for collaborative projects. Out of these eight requests, three were completed and one was deemed unfeasible. User requests cover a wide range of domains in biology and image analysis (Table 2), which highlights the diversity of expertise F-BIAS can offer, with a distributed team coming from diverse backgrounds.
@@
Users can request a collaborative project by direct contact, via Euro-BioImaging user access website or during an open desk session. The project registration form developed by Euro-BioImaging provides a template for users to specify the project content. However due to the diversity of the projects handled by F-BIAS (which can, for instance, extend beyond light microscopy), it is challenging to create a universal template that perfectly captures all requests while remaining concise. As a result, we always discuss with the user before making any commitment. The request is then discussed among the F-BIAS team - that serves as a selection committee - to evaluate the project’s feasibility. If deemed feasible, one analyst with the required expertise and availability takes over the project. It evolves iteratively with regular discussions between the biologist and the analyst. Additionally, the project’s analyst may be paired with another analyst from a different site providing technical mentorship. This again results in capitalizing on user services to improve the skills and expertise of analysts (Figure 2B). Given the current overbooked status of analysts, projects are allocated solely based on their expertise and availability, and potential outcomes have only little impact in the decision process.
Unlike consultations, collaborative projects are billed. We use an hourly fee to charge the user for the projected effort of the project. When the actual effort exceeds the planned duration by a significant amount, a new quote is prepared in agreement with the user. We leveraged our virtual and decentralized organization to establish a simplified economic model for F-BIAS. In the inception phase described here, we rely on the existing recharge systems implemented by contributing core facilities. The quote for a project handled by an F-BIAS analyst, funded centrally by FBI, is sent from the local core facility in which they are embedded, and the fees are received by the core facility. In turn, the core facility hosts and manages the analyst, handling all administrative overhead. The hourly fees are subsidized by FBI to lower and standardize the project cost to the user, ensuring that all F-BIAS analysts bill the same amount regardless of the core facility they work in. This subsidy can take the form of funds or in-kind contributions, with the F-BIAS analyst funded by FBI dedicating part of their time to work on user projects for the local core. This organizational structure is straightforward and has the advantage of requiring minimal effort to set up. This demonstrates the commitment and importance of a national infrastructure like FBI, which can leverage funds to support the dissemination of bioimage analysis services at the national scale.
@@
Figure 2: Overview of the bioimage analysis services provided by F-BIAS. A. The consultations services take the shape of open desk sessions of one hour organized regularly. A user can book such a session in an online form, providing a description of the image analysis problem they want to discuss, and uploading a few example images. During the preparation phase, the requests are each allocated to an open desk session, with a panel of 2 to 3 analysts selected according to their expertise. All the sessions happen the same day, with 2 to 3 running in parallel. The analysts join the user in an online meeting of 1 hour and work on the user’s problem during this time, actively prototyping the solutions they propose. The output of a session can be advice on how to address the problem, small analysis pipeline or small macros. B. When a question cannot be addressed in one hour and / or requires some development, a user can request a collaborative project, during an open desk session or by direct contact. Project requests are then evaluated by the F-BIAS team, which estimates the project cost, start delay, and duration. Upon acceptance, an analyst is assigned based on availability and expertise. As a potential opportunity. a senior analyst from another site will provide mentoring to the project’s analyst. The user and analyst collaborate asynchronously, with regular analyst reports, using various communication media (emails, online/onsite meetings). Project outputs may include analysis pipelines, results, and reports.
@@
The added value of a distributed image analysis core facility
@@
F-BIAS successfully addresses isolation issues mentioned in the above. Within F-BIAS, analysts exchange knowledge and good practices, teach each other analysis techniques and share information on key publications, online resources or conferences. Moreover, effective communication about the ongoing projects helps prevent duplication of work on similar project, saving both time and resources. Experienced members can also mentor early-career peers and help them navigate the global bioimage analysis world. Conversely, early-career analysts challenge the techniques senior analysts are used to relying on, and help them renew their technical portfolio. This is particularly valuable given the young age of the field in academia. The consultations with several analysts amounted to an opportunity for learning how others approach a new problem and exchanging ideas. It also provided the group’s developers with a chance to demonstrate new tools and developments, have them tested by other analysts, and receive feedback. F-BIAS is highly valued by its members as a vital platform for honing skills, exploring new techniques, and discovering new tools.
@@
The scientific community of France is the main target of this endeavor, especially biologists who do not have access to bioimage analysis expertise onsite, the most common situation in the country. These scientists now have access to the expertise of about 20 analysts and the means to leverage it for their projects. Consultations are tailored for efficiency and maximizing the scientific output of a one-hour session. F-BIAS analysts master a wide range of image analysis languages and packages from scripting to advanced analysis modules and AI libraries. By describing their needs in the online registration form, users will be put in contact with analysts that have the given expertise to answer their question, when possible. From the user viewpoint, this process is much simpler than having to identify the required expertise to find the right facility to contact. This one-hour consulting service with several experts is free of charge and requires little effort beside registration and sending test data beforehand. Finally, when a question cannot be solved in one hour and the user does not have resources to develop a suggested solution, F-BIAS can initiate a collaborative project with the user. F-BIAS truly succeeds in overcoming the technological barriers associated with bioimage analysis for researchers within France.
@@
The integration of F-BIAS within France-BioImaging enables it to offer an efficient service in image analysis that complements other services. Moreover, F-BIAS reached critical size to foster beneficial interactions with other projects of France-BioImaging related to bioimage analysis. For instance, the FBI.Data project is focused on scientific image data management, offering tools for intelligent data archival, high performance computing, and implementation of FAIR data principles. A federated data management solution (29) is currently being deployed in the different nodes of the FBI infrastructure, in strong connection with shared computing resources. FBI.Data can capitalize on F-BIAS to test and deploy its solutions in host core facilities, and reach out to end users, using F-BIAS members as a medium. FBI also supports the development of large scientific framework for image analysis such as Icy (30), BioImageIT (31) and MorphoNet (32). F-BIAS members serve as early adopters, beta testers and feedback source for these projects, offering a wide range of biological applications to disseminate the software through the F-BIAS user base. Finally, F-BIAS can redirect projects that fall outside the scope of its services to research groups of FBI focused on image processing.
For industry partners
F-BIAS is in the best position to recommend an expert in image analysis to industrial companies, who can address their challenges. For now, this type of request is handled locally by the analyst’s core facility in collaboration with the company.
@@
The most significant challenge is to secure human resources, in particular for the collaborative project services. A major issue in operating the group and bioimage analysis core facilities in general is the limited number of dedicated analysts compared to the needs of the biological community. Indeed, the funding from the institutions to secure analyst positions dedicated to service are too restricted to have a sufficient workforce. F-BIAS is a virtual structure, initially built on existing resources that are federated from host core facilities. At its creation, all members were already saturated or close to being saturated with requests from their local scientific community. This low availability is the main factor delaying image analysis projects (18). The inception of F-BIAS was facilitated by the scientific value it offered to members and the minimal time commitment required for participation in consultations and exchanges. This helped in building and maintaining a sufficiently large team of willing analysts. Additionally, the positions granted by FBI for collaborative projects further strengthened our services.
@@
Ensuring the scientific added value of internal meetings
The scientific added value analysts find in F-BIAS is crucial for the success of the initiative. It helped convince the analysts’ managers to allocate time to the initiative, and helped retain analysts and prevent dropouts. It is therefore important to control the content of F-BIAS regular and annual meetings. We do this by enforcing the rules described above and focus on technical exchanges. Regular meetings are mainly focused on scientific content, asking for each other’s help or showing a new technique, and round tables on more organizational aspects are reserved for our annual hackathons.
@@
The desirable components of a distributed image analysis core facility
There are a few other initiatives building nationwide services for bioimage analysis. Among them we can mention the SciLifeLab BioImage Informatics infrastructure in Sweden (34) and the National Facility for Data Handling & Analysis in Italy (35). They have the advantage of enjoying human resources dedicated to these projects from their inception. SciLifeLab has units at various sites (not all analysts are based at the same university), with agreements in place to govern personnel responsibilities, payments, and other related matters. On our side we started as a virtual structure, built from existing resources initially with other duties, spread over the country. Our approach described above managed to circumvent these challenges and even converted them into opportunities. Nonetheless we identified a few key components, additional to the human resources and online tools discussed above, that would facilitate reproducing such a service elsewhere. As we did not benefit from all of them immediately, we can attest to their importance.